OpenStudy (anonymous):

The first four terms of a sequence are shown below: 6, 2, -2, -6 Which of the following functions best defines this sequence?

OpenStudy (anonymous):

f(1) = 6, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 2; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) + 2; for n ≥ 1

Mehek (mehek14):

how much do you add or subtract to 6 to get 2?

OpenStudy (anonymous):

4

Mehek (mehek14):

add or subtract?

OpenStudy (anonymous):

subtract

Mehek (mehek14):

so -4

Mehek (mehek14):

how about from 2 to -2?

Mehek (mehek14):

and -2 to -6?

OpenStudy (anonymous):

subtract 4

Mehek (mehek14):

so f(1) = 6, f(n + 1) = f(n) - 4; for n ≥ 1

OpenStudy (anonymous):

thanks

Mehek (mehek14):

yw